OpenStudy (anonymous):

(7xy^3+5x^3y^4)-(4xy^3-2x^3y^4) Subtract the polynomials.

OpenStudy (anonymous):

first you find like terms in the first and second sets of parenthesis and use those to do simple subtraction

Directrix (directrix):

Do you know the Distributive Property? @reason (7xy^3+5x^3y^4)-(4xy^3-2x^3y^4)= 7xy^3 + 5x^3y^4) -1* (4xy^3-2x^3y^4)=

OpenStudy (anonymous):

I know you have to change the sign to plus and the second part becomes plus -4xy^3+2x....etc

Directrix (directrix):

( 7xy^3 + 5x^3y^4) -1* ( 4xy^3 -2x^3y^4 )= ( 7xy^3 + 5x^3y^4) + (-1)*( 4xy^3 ) + (-1) (-2x^3y^4) = So the negative one is distributed over the two terms in the second expression. ( 7xy^3 + 5x^3y^4) + (-1)*( 4xy^3 ) + (-1) (-2x^3y^4) = 7xy^3 + 5x^3y^4 - 4xy^3 + ____?_____ What goes here? @reason

Directrix (directrix):

@ reason Look at just this part: (-1) (-2x^3y^4) --> What is negative one times negative 2 x cubed y to the fourth?

OpenStudy (anonymous):

i'm having trouble with this one

OpenStudy (anonymous):

=(

Directrix (directrix):

(-1) (-2x^3y^4) = 2x^3y^4

Directrix (directrix):

( 7xy^3 + 5x^3y^4) + (-1)*( 4xy^3 ) + (-1) (-2x^3y^4) = 7 xy^3 + 5 x^3y^4 - 4 xy^3 + 2 x^3y^4 = 3 x y^3 + 7 x^3 y -->@reason